If 20℅ of A= B & 40℅ of B=C then 60℅ of (A+B)=?
Answer by math_helper(2461)

0.2A = B —> A = 5B
0.4B = C —> B = 2.5C
I assume the problem wants the answer in terms of C:
--
(0.6)(A+B) =
(0.6)(5B + 2.5C) = (substituted for A and B)
(0.6)(5(2.5C) + 2.5C) = (substituted for B)
(0.6)(15C) =
9C